### v1

# arXiv:2602.23464v1[cs.CR] 26 Feb 2026

Abstract
Multi-scalar multiplication (MSM), $\mathsf{MSM}(\vec{P},\vec{x})=\sum_{i=1}^{n}x_{i}P_{i}$, is a dominant computational kernel in discrete-logarithm–based cryptography and often becomes a bottleneck for verifiers and other resource-constrained clients. We present 2G2T, a simple protocol for verifiably outsourcing MSM to an untrusted server. After a one-time keyed setup for fixed bases $\vec{P}$ (producing a public merged-bases vector $\vec{T}$ and client secret state), the server returns only two group elements per query: the claimed MSM result $A=\mathsf{MSM}(\vec{P},\vec{x})$ and an auxiliary value $B=\mathsf{MSM}(\vec{T},\vec{x})$. Verification requires a single length-$n$ field inner product and a constant number of group operations (two scalar multiplications and one addition), while the server performs two MSMs. In our Ristretto255 implementation, verification is up to $\sim 300\times$ faster than computing the MSM locally using a highly optimized MSM routine for $n$ up to $2^{18}$, and the server-to-client response is constant-size (two compressed group elements, 64 bytes on Ristretto255). Despite its simplicity and efficiency, 2G2T achieves statistical soundness: for any (even unbounded) adversarial server, the probability of accepting an incorrect result is at most $1/q$ per query, and at most $e/q$ over $e$ adaptive executions, in a prime-order group of size $q$.
1 Introduction
Multi-scalar multiplication (MSM) is a ubiquitous computational kernel in prime-order groups and lies at the heart of modern discrete-logarithm–based cryptography. Let $\mathbb{G}$ be a cyclic group of prime order $q$, written additively, with scalar field $\mathbb{F}_{q}$. Given bases $\vec{P}=(P_{1},\dots,P_{n})\in\mathbb{G}^{n}$ and scalars $\vec{x}=(x_{1},\dots,x_{n})\in\mathbb{F}_{q}^{n}$, the MSM is
$\mathsf{MSM}(\vec{P},\vec{x})\;:=\;\sum_{i=1}^{n}x_{i}P_{i}\;\in\;\mathbb{G}.$
Even with state-of-the-art algorithms (e.g., Pippenger-style methods), MSM computation requires many group operations and remains costly for large $n$ (see, e.g., [2]).
A key reason MSM is so prominent is that many protocols deliberately aggregate many checks into a small number of large group equations, so verification is dominated by one (or a small number of) large MSMs. In our target applications, these MSMs are not only large, but also fixed-base across many queries: a system publishes (or deterministically derives) a long vector of public generators once, and then many subsequent proofs/commitments are verified by taking different linear combinations of the same bases. Bulletproofs is a canonical example: verification of a range proof—and especially aggregated verification—boils down to a large MSM in public generator vectors, and the protocol supports aggregation and batching to reduce marginal verification cost [4].
A second major source of large, fixed-base MSMs is vector commitments and inner-product–based arguments. In IPA-based polynomial commitments as used in Halo/Halo2-style systems, the commitment key contains a vector $\vec{G}\in\mathbb{G}^{d}$ of $d$ group elements, and committing evaluates an inner product $\langle\vec{a},\vec{G}\rangle$, i.e., an MSM over $d$ bases [20, 3]. The Halo2 documentation makes explicit that $d=2^{k}$; consequently, even moderate parameters (e.g., $k=11$ so $d=2048$) already induce MSMs with thousands of terms, while larger $k$ scales $d$ exponentially [20, 19]. In deployments, such long generator vectors are typically instantiated in a reproducible “nothing-up-my-sleeve” manner—for example by hashing domain-separated strings to curve points using standardized hash-to-curve techniques [7]—and then reused across many verifications.
Beyond zero-knowledge, related fixed-base multi-exponentiation/MSM patterns also arise in distributed key generation (DKG) [16, 10] and verifiable secret sharing (VSS) [8, 17]: parties verify share-consistency relations against a fixed vector of public commitments (typically coefficient commitments), and these checks reduce to multi-exponentiations/MSMs over those fixed bases.
Outsourcing motivation. Outsourcing becomes attractive when the MSM dimension $n$ is large enough that computing $\mathsf{MSM}(\vec{P},\vec{x})$ locally dominates cost under latency, energy, or throughput constraints. This arises in two common scenarios: (i) resource-constrained clients/verifiers (mobile/embedded devices, thin clients, or applications running under tight energy/latency budgets) that must validate many proofs or commitments but cannot afford repeated large MSMs; and (ii) high-throughput verifiers that validate large volumes of statements, where MSM lies on the critical path and must be aggressively parallelized to keep up with demand. In both settings, the fastest MSM implementations rely on careful engineering and often benefit from parallel hardware (e.g., wide SIMD and, in some deployments, GPUs/accelerators) that may be unavailable or impractical to deploy at the verifier, motivating delegation to a more powerful server. Finally, in some deployments the natural server is already present: the proof generator/prover already holds the coefficient vector $\vec{x}$ and can compute the MSM on behalf of the verifier.
Delegation, however, immediately introduces an integrity problem: an untrusted server may be faulty or adversarial and can return an arbitrary group element. Thus, MSM outsourcing falls squarely within the verifiable computation setting: the client seeks strong correctness guarantees while expending substantially less work than locally computing the MSM. This tension is formalized in foundational work on verifiable delegation and verifiable computing [12, 9]. General-purpose verifiable computation is powerful but often too heavy to deploy at the granularity of a ubiquitous cryptographic kernel. This motivates specialized protocols tailored to MSM that provide strong integrity guarantees with minimal client-side verification overhead.
Challenges. Prior work on verifiable MSM outsourcing highlights two obstacles: (i) achieving a meaningful speedup while keeping statistical soundness extremely high, and (ii) avoiding lightweight checks that enforce only a consistency relation between returned values but do not bind the server to all $n$ terms of the instance. For example, Pinkas [18] gives a non-interactive, publicly verifiable statistical check that can yield a modest constant-factor speedup when one targets moderate soundness (on the order of tens of bits), but doing so typically increases the server’s response size to hundreds of group elements on 255-bit groups, and the speedup diminishes as the soundness target approaches “full-group” levels. At the other end of the spectrum, cheaper consistency-only strategies can fail outright against a malicious server: a server may compute an MSM over only a strict subset of indices (e.g., return $A=x_{1}P_{1}$) and still satisfy the verifier’s consistency relation, even though $A\neq\mathsf{MSM}(\vec{P},\vec{x})$ (see, e.g., [21]). Together, these examples underscore that verifiable MSM outsourcing is nontrivial: one must simultaneously achieve negligible soundness error and bind the server to the full MSM instance without reintroducing MSM-like verifier work.
Our contribution: 2G2T. We present 2G2T, a statistically sound protocol for verifiable MSM outsourcing with constant-size server-to-client communication and fast client-side verification. [^1] The mnemonic 2G2T reflects (i) “two group elements to transfer”: the server returns two group elements per query; and (ii) it can seem “too good to be true” at first glance, as it combines constant-size communication, $O(1)$ verifier group operations, and $1/q$-level statistical soundness. After a one-time setup for fixed bases $\vec{P}$, the server answers each client’s query $\vec{x}$ with only two group elements: $A$ claimed to equal $\mathsf{MSM}(\vec{P},\vec{x})$ and an auxiliary value $B=\mathsf{MSM}(\vec{T},\vec{x})$ for a precomputed merged-bases vector $\vec{T}$. The client verifies using a single length-$n$ field inner product $s=\langle\vec{x},\vec{\rho}\rangle$ and a constant number of group operations (two scalar multiplications and one addition), while the server performs exactly two MSMs (only one more than computing $A$ alone). We prove that 2G2T is perfectly complete and statistically sound: for any (even computationally unbounded) adversarial server, the probability that the verifier accepts an incorrect $A$ is at most $1/q$ per execution, and at most $e/q$ over $e$ adaptive executions (Theorems 4.14.3).
Our Rust implementation over Ristretto255 confirms that 2G2T yields large practical gains in the large-$n$ regime: verification is up to $\sim 300\times$ faster than computing a highly optimized MSM, and up to $\sim 3{,}000\times$ faster than a naïve “$n$ scalar multiplications” baseline, for dimensions up to $n=2^{18}=262{,}144$ (with exact constants depending on hardware and implementation details). Finally, server-to-client communication is constant and minimal: the response consists of two compressed group elements $(A,B)$, i.e., 64 bytes in Ristretto255, independent of $n$. These efficiency gains come while providing “full-group” statistical integrity: the per-query soundness error $1/q$ corresponds to about $\log_{2}(q)$ bits of soundness (roughly $252$ bits for Ristretto255).
2 Threat model and security goals
We consider the verifiable outsourcing setting for MSM. A (potentially resource-constrained) client $\mathcal{C}$ holds bases $\vec{P}\in\mathbb{G}^{n}$ and, for each query, coefficients $\vec{x}\in\mathbb{F}_{q}^{n}$, and wishes to obtain $\mathsf{MSM}(\vec{P},\vec{x})$ with substantially less online work than computing the MSM locally. The client interacts with an external server $\mathcal{S}$ that is powerful but untrusted.
Adversarial server.
The server may deviate arbitrarily from the prescribed algorithm: it may be faulty, malicious, or economically motivated to return an incorrect result to save computation. We model $\mathcal{S}$ as an adaptive adversary that, after observing the public setup information and the full transcript of all prior executions (including whether the client accepted or rejected), can choose future queries $\vec{x}$ and responses as an arbitrary function of its entire view. We allow $\mathcal{S}$ to be computationally unbounded; thus our integrity guarantees are information-theoretic/statistical rather than based on computational hardness.
Communication model.
We assume the network provides authenticity of messages (or equivalently, that the client can bind responses to the corresponding request and setup state); confidentiality of network traffic is not required for soundness.
Information revealed.
We do not aim to hide the scalars $\vec{x}$ from the server (this corresponds to the common verifier-side use case where $\vec{x}$ is public, or already known to the party doing the computation). Similarly, we do not require privacy for the MSM output beyond what is implied by the application. The client keeps its protocol-specific secret state (e.g., the key material produced in the setup phase) confidential.
Security goals.
An MSM outsourcing protocol should satisfy the following standard properties.
  • Completeness. If $\mathcal{S}$ follows the protocol and returns correct values, then $\mathcal{C}$ accepts and outputs $Y=\mathsf{MSM}(\vec{P},\vec{x})$ for every query $\vec{x}$.
  • Soundness (integrity). No adversarial server $\mathcal{S}^{\ast}$ should be able to make $\mathcal{C}$ accept an incorrect MSM result except with small probability. In our setting, we target statistical soundness: for any (even computationally unbounded) $\mathcal{S}^{\ast}$ and any adaptively chosen queries, the probability that the client accepts while outputting $A\neq\mathsf{MSM}(\vec{P},\vec{x})$ is negligible in the chosen soundness/security parameter (often written as $\leq 2^{-\lambda}$ for $\lambda$-bit statistical soundness). In 2G2T, this probability is at most $1/q$ per query (equivalently, about $\log_{2}(q)$ bits).
  • Efficiency. The client’s online work per query (including verification) should be significantly cheaper than locally computing $\mathsf{MSM}(\vec{P},\vec{x})$. We also seek small communication overhead.
3 The 2G2T Protocol
At a high level, 2G2T rests on two basic observations that become powerful when combined. First, MSM is linear in the bases (under coordinate-wise operations): for all $\vec{P},\vec{P}^{\prime}\in\mathbb{G}^{n}$, $\alpha\in\mathbb{F}_{q}$, and $\vec{x}\in\mathbb{F}_{q}^{n}$,
$\mathsf{MSM}(\vec{P}+\vec{P}^{\prime},\vec{x})=\mathsf{MSM}(\vec{P},\vec{x})+\mathsf{MSM}(\vec{P}^{\prime},\vec{x})\quad\text{and}\quad\mathsf{MSM}(\alpha\vec{P},\vec{x})=\alpha\mathsf{MSM}(\vec{P},\vec{x}),$
where $\vec{P}+\vec{P}^{\prime}$ and $\alpha\vec{P}$ denote coordinate-wise group addition and scalar multiplication. Second, MSM becomes particularly cheap when all bases lie in the span of a single point $Q\in\mathbb{G}$ and the verifier knows the corresponding coefficients: for $\vec{Q}=(\rho_{1}Q,\dots,\rho_{n}Q)$ with known $\vec{\rho}=(\rho_{1},\dots,\rho_{n})\in\mathbb{F}_{q}^{n}$,
$\mathsf{MSM}(\vec{Q},\vec{x})\;=\;\sum_{i=1}^{n}x_{i}(\rho_{i}Q)\;=\;\langle\vec{x},\vec{\rho}\rangle\,Q,$
which reduces to a field inner product and a single fixed-base scalar multiplication by $Q$.
2G2T leverages the above facts by having the client precompute a merged-bases vector
$\vec{T}\;=\;r\vec{P}+(\rho_{i}Q)_{i}\;=\;(rP_{1}+\rho_{1}Q,\dots,rP_{n}+\rho_{n}Q),$
for a secret scalar $r\in\mathbb{F}_{q}$ and secret vector $\vec{\rho}\in\mathbb{F}_{q}^{n}$, and provide $\vec{T}$ to the server (which may cache it for subsequent queries). On input a query vector $\vec{x}$, the server returns
$A\stackrel{{\scriptstyle?}}{{=}}\mathsf{MSM}(\vec{P},\vec{x})\qquad\text{and}\qquad B\stackrel{{\scriptstyle?}}{{=}}\mathsf{MSM}(\vec{T},\vec{x}).$
The client verifies using the relation $B\stackrel{{\scriptstyle?}}{{=}}rA+\langle\vec{x},\vec{\rho}\rangle Q$. Intuitively, acceptance forces the server’s response to be consistent with a single hidden scalar $r$ and a verifier-known fixed-base component, yielding statistical soundness while keeping the verifier far below the cost of an MSM. We give the full protocol description below, and summarize it in pseudocode in Algorithm 1.
Precomputation (once).
Given public bases $\vec{P}=(P_{1},\dots,P_{n})\in\mathbb{G}^{n}$ and a fixed public non-identity point $Q\in\mathbb{G}\setminus\{0\}$, the client samples a secret scalar $r\xleftarrow{\mathdollar}\mathbb{F}_{q}^{\*}$ and a secret random vector $\vec{\rho}=(\rho_{1},\dots,\rho_{n})\xleftarrow{\mathdollar}\mathbb{F}_{q}^{n}$. It defines the (implicit) trap bases
$\vec{Q}:=(\rho_{1}Q,\dots,\rho_{n}Q)\in\mathbb{G}^{n}$
and computes the merged-bases vector
$\vec{T}\;:=\;r\vec{P}+\vec{Q}\;=\;(rP_{1}+\rho_{1}Q,\;\dots,\;rP_{n}+\rho_{n}Q)\in\mathbb{G}^{n}.$
The client uploads $\vec{T}$ to the server (and $\vec{P}$ if not already known), and may treat $\vec{T}$ as a long-lived public key while keeping $(r,\vec{\rho})$ as a long-lived secret key.
MSM outsourcing (per query).
To outsource the computation of $\mathsf{MSM}(\vec{P},\vec{x})$, the client sends $\vec{x}$ to the server. (If the server has not cached $\vec{P}$ and/or $\vec{T}$, the client sends them as well.) The server returns two group elements
$A\stackrel{{\scriptstyle?}}{{=}}\mathsf{MSM}(\vec{P},\vec{x})\qquad\text{and}\qquad B\stackrel{{\scriptstyle?}}{{=}}\mathsf{MSM}(\vec{T},\vec{x}).$
Verification.
The client computes the inner product
$s\;:=\;\langle\vec{x},\vec{\rho}\rangle\;=\;\sum_{i=1}^{n}x_{i}\rho_{i}\in\mathbb{F}_{q}$
and accepts iff
$B\;\stackrel{{\scriptstyle?}}{{=}}\;rA+sQ.$
Algorithm 1 2G2T: verifiable outsourcing for $\mathsf{MSM}(\vec{P},\vec{x})$
1:Input: public $\mathbb{G}$ of prime order $q$, public $Q\in\mathbb{G}\setminus\{0\}$, bases $\vec{P}=(P_{1},\dots,P_{n})\in\mathbb{G}^{n}$
2:Input (per query): coefficients $\vec{x}=(x_{1},\dots,x_{n})\in\mathbb{F}_{q}^{n}$
3:Output (per query): $A\in\mathbb{G}$ or $\bot$
4:
5:Precomputation / key setup (once).
6:Sample $r\xleftarrow{\mathdollar}\mathbb{F}_{q}^{\*}$ and $\vec{\rho}=(\rho_{1},\dots,\rho_{n})\xleftarrow{\mathdollar}\mathbb{F}_{q}^{n}$
7:for $i\leftarrow 1$ to $n$ do
8:  $T_{i}\leftarrow rP_{i}+\rho_{i}Q$
9:end for
[…]
11:Upload $\vec{T}$ to the server (and $\vec{P}$ if needed) and have the server cache them $\triangleright$ Client keeps $\mathsf{sk}:=(r,\vec{\rho})$ secret
12:
13:Online outsourcing (per query for $\vec{x}$).
14:Client $\to$ Server: send $\vec{x}$
15:Server: compute $A\leftarrow\mathsf{MSM}(\vec{P},\vec{x})$ and $B\leftarrow\mathsf{MSM}(\vec{T},\vec{x})$
16:Server $\to$ Client: return $(A,B)$
17:
18:Client verification (per query).
19:Compute $s\leftarrow\langle\vec{x},\vec{\rho}\rangle=\sum_{i=1}^{n}x_{i}\rho_{i}\in\mathbb{F}_{q}$
20:if $B\neq rA+sQ$ then
21:  return $\bot$
22:end if
23:return AA
4 2G2T: Completeness and Soundness
Common setting.
Throughout the remainder of this section, we work in the following setup.
Definition 1(2G2T setup)
Let $\mathbb{G}$ be a cyclic group of prime order $q$ written additively, with scalar field $\mathbb{F}_{q}$. Let $Q\in\mathbb{G}\setminus\{0\}$ be public and fix $\vec{P}=(P_{1},\dots,P_{n})\in\mathbb{G}^{n}$. The 2G2T precomputation samples $r\xleftarrow{\mathdollar}\mathbb{F}_{q}$ and $\vec{\rho}=(\rho_{1},\dots,\rho_{n})\xleftarrow{\mathdollar}\mathbb{F}_{q}^{n}$, and defines $\vec{T}\in\mathbb{G}^{n}$ by
$T_{i}:=rP_{i}+\rho_{i}Q\qquad\text{for all }i\in[n].$
Given a query $\vec{x}\in\mathbb{F}_{q}^{n}$ and a server response $(A,B)\in\mathbb{G}^{2}$, the verifier runs the check in Algorithm 1 and we denote its decision by $\mathsf{Vrfy}(r,\vec{\rho};\vec{x},A,B)\in\{0,1\}$.
Theorem 4.1(Perfect completeness of 2G2T)
In the setting of Definition 1, for every $\vec{x}\in\mathbb{F}_{q}^{n}$, if the server returns
$A=\mathsf{MSM}(\vec{P},\vec{x})\qquad\text{and}\qquad B=\mathsf{MSM}(\vec{T},\vec{x}),$
then the verifier in Algorithm 1 accepts and outputs AA.
Proof
Let $s:=\langle\vec{x},\vec{\rho}\rangle=\sum_{i=1}^{n}x_{i}\rho_{i}\in\mathbb{F}_{q}$. Using $T_{i}=rP_{i}+\rho_{i}Q$ and expanding the MSM,
$B\;=\;\sum_{i=1}^{n}x_{i}T_{i}\;=\;\sum_{i=1}^{n}x_{i}(rP_{i}+\rho_{i}Q)\;=\;r\sum_{i=1}^{n}x_{i}P_{i}\;+\;\left(\sum_{i=1}^{n}x_{i}\rho_{i}\right)Q\;=\;rA+sQ.$
Hence the verifier’s check $B\stackrel{{\scriptstyle?}}{{=}}rA+sQ$ passes and it outputs AA.∎
Next, we present two lemmas that will be used in the soundness proofs (Theorems 4.2 and 4.3).
Lemma 1(Independence of $r$ and $\vec{T}$)
In the setting of Definition 1, the random variables $r$ and $\vec{T}$ are statistically independent. Equivalently, for every $t\in\mathbb{G}^{n}$ and every $a\in\mathbb{F}_{q}$,
$\Pr[r=a\mid\vec{T}=t]\;=\;\frac{1}{q}.$
Proof
Fix any $a\in\mathbb{F}_{q}$ and condition on $r=a$. Then for each $i\in[n]$,
$T_{i}\;=\;aP_{i}+\rho_{i}Q,\qquad\rho_{i}\xleftarrow{\mathdollar}\mathbb{F}_{q}\text{ independently.}$
Since $\mathbb{G}$ has prime order and $Q\neq 0$, the map $\rho\mapsto\rho Q$ is a bijection $\mathbb{F}_{q}\to\mathbb{G}$, so $\rho_{i}Q$ is uniform in $\mathbb{G}$. Adding the fixed point $aP_{i}$ is a translation, hence $T_{i}$ is uniform in $\mathbb{G}$. Independence of the $\rho_{i}$ implies that $\vec{T}$ is uniform in $\mathbb{G}^{n}$, i.e., for every $t\in\mathbb{G}^{n}$,
$\Pr[\vec{T}=t\mid r=a]\;=\;\frac{1}{q^{n}},$
which does not depend on $a$. Therefore $\Pr[\vec{T}=t\mid r=a]=\Pr[\vec{T}=t]$ for all $a,t$, thus $r$ and $\vec{T}$ are independent. Since $r\xleftarrow{\mathdollar}\mathbb{F}_{q}$, we have $\Pr[r=a]=1/q$, and by independence
$\Pr[r=a\mid\vec{T}=t]\;=\;\Pr[r=a]\;=\;\frac{1}{q}.$
Lemma 2(Uniqueness of an accepting scalar for an incorrect $A$)
Fix $\vec{P}\in\mathbb{G}^{n}$, $\vec{T}\in\mathbb{G}^{n}$, a query $\vec{x}\in\mathbb{F}_{q}^{n}$, and a pair $(A,B)\in\mathbb{G}^{2}$. Let $A^{\star}:=\mathsf{MSM}(\vec{P},\vec{x})$ and for each $r\in\mathbb{F}_{q}$ define $\vec{Q}^{(r)}:=\vec{T}-r\vec{P}\in\mathbb{G}^{n}$. If there exist two distinct scalars $r_{1}\neq r_{2}$ such that
$B\;=\;r_{1}A+\mathsf{MSM}(\vec{Q}^{(r_{1})},\vec{x})\qquad\text{and}\qquad B\;=\;r_{2}A+\mathsf{MSM}(\vec{Q}^{(r_{2})},\vec{x}),$
then $A=A^{\star}$. In particular, if $A\neq A^{\star}$, then there is at most one scalar $r\in\mathbb{F}_{q}$ for which the verifier’s check can accept $(A,B)$.
Proof
Subtract the two equalities:
$0\;=\;(r_{2}-r_{1})A+\mathsf{MSM}(\vec{Q}^{(r_{2})}-\vec{Q}^{(r_{1})},\vec{x}).$
Since $\vec{Q}^{(r_{2})}-\vec{Q}^{(r_{1})}=(\vec{T}-r_{2}\vec{P})-(\vec{T}-r_{1}\vec{P})=(r_{1}-r_{2})\vec{P}$, linearity of MSM gives
$\mathsf{MSM}(\vec{Q}^{(r_{2})}-\vec{Q}^{(r_{1})},\vec{x})\;=\;(r_{1}-r_{2})\mathsf{MSM}(\vec{P},\vec{x})\;=\;(r_{1}-r_{2})A^{\star}.$
Hence
$0\;=\;(r_{2}-r_{1})A+(r_{1}-r_{2})A^{\star}\;=\;(r_{2}-r_{1})(A-A^{\star}).$
Because $r_{2}\neq r_{1}$, we have $r_{2}-r_{1}\neq 0$ in $\mathbb{F}_{q}$, thus $A=A^{\star}$.∎
Theorem 4.2(Statistical soundness of 2G2T)
In the setting of Definition 1, for any (possibly computationally unbounded) adversarial server $\mathcal{S}^{\*}$ that, on input $(\vec{P},\vec{T},\vec{x})$ for an arbitrary query $\vec{x}\in\mathbb{F}_{q}^{n}$, outputs $(A,B)\in\mathbb{G}^{2}$, we have
$\Pr\Bigl[\,\mathsf{Vrfy}(r,\vec{\rho};\vec{x},A,B)=1\ \wedge\ A\neq\mathsf{MSM}(\vec{P},\vec{x})\,\Bigr]\;\leq\;\frac{1}{q},$
where the probability is over the client’s randomness in the precomputation of Definition 1.
Proof
Fix an arbitrary query $\vec{x}\in\mathbb{F}_{q}^{n}$ and let $(A,B)\in\mathbb{G}^{2}$ be the output of $\mathcal{S}^{\*}$ on input $(\vec{P},\vec{T},\vec{x})$. Write $A^{\star}:=\mathsf{MSM}(\vec{P},\vec{x})$.
Fix an arbitrary $t\in\mathbb{G}^{n}$ and condition on the event $\vec{T}=t$ (and on $\mathcal{S}^{\*}$’s internal randomness, so that $(A,B)$ is fixed). By Lemma 1, the secret scalar $r$ is uniform over $\mathbb{F}_{q}$ under this conditioning.
For each $\hat{r}\in\mathbb{F}_{q}$, define the derived vector
$\vec{Q}^{(\hat{r})}\;:=\;\vec{T}-\hat{r}\,\vec{P}\;=\;t-\hat{r}\,\vec{P}\in\mathbb{G}^{n}.$
In particular, for the actual secret $r$ used in setup, the relation $\vec{T}=r\vec{P}+(\rho_{i}Q)_{i}$ implies $\vec{Q}^{(r)}=(\rho_{i}Q)_{i}$, and hence
$\mathsf{MSM}(\vec{Q}^{(r)},\vec{x})\;=\;\mathsf{MSM}\bigl((\rho_{i}Q)_{i},\vec{x}\bigr)\;=\;\langle\vec{x},\vec{\rho}\rangle Q.$
Therefore the verifier’s check $B\stackrel{{\scriptstyle?}}{{=}}rA+\langle\vec{x},\vec{\rho}\rangle Q$ is equivalent to
$B\;=\;rA\;+\;\mathsf{MSM}(\vec{Q}^{(r)},\vec{x}).$
If $A\neq A^{\star}$, then by Lemma 2 the above equality can hold for at most one value of $r\in\mathbb{F}_{q}$. Since $r$ is uniform in $\mathbb{F}_{q}$ conditioned on $\vec{T}=t$, it follows that
$\Pr\Bigl[\mathsf{Vrfy}(r,\vec{\rho};\vec{x},A,B)=1\ \wedge\ A\neq A^{\star}\ \Bigm|\ \vec{T}=t\Bigr]\;\leq\;\frac{1}{q}.$
Since $t$ was arbitrary, the same bound holds without conditioning, proving the theorem. ∎
Theorem 4.3($e$-execution statistical soundness of 2G2T)
In the setting of Definition 1, let $\mathcal{S}^{\*}$ be any (possibly computationally unbounded) adaptive adversary that is given $(\vec{P},\vec{T})$ and then interacts with the verifier for at most $e$ executions. In execution $j\in[e]$, $\mathcal{S}^{\*}$ adaptively chooses a query $\vec{x}^{(j)}\in\mathbb{F}_{q}^{n}$ and outputs a pair $(A^{(j)},B^{(j)})\in\mathbb{G}^{2}$, after which the verifier returns the accept/reject bit. Then
$\Pr\Bigl[\exists\,j\in[e]\!:\ \mathsf{Vrfy}(r,\vec{\rho};\vec{x}^{(j)},A^{(j)},B^{(j)})=1\ \wedge\ A^{(j)}\neq\mathsf{MSM}(\vec{P},\vec{x}^{(j)})\Bigr]\;\leq\;\frac{e}{q},$
where the probability is over the client’s randomness in the precomputation and the internal randomness of $\mathcal{S}^{\*}$.
Proof
Fix any adaptive adversary $\mathcal{S}^{\*}$ and consider the interaction for at most $e$ executions. For $j\in\{1,\dots,e\}$, let $\mathsf{Bad}_{\leq j}$ be the event that an incorrect acceptance occurs in one of the first $j$ executions, and let $\mathsf{Bad}:=\mathsf{Bad}_{\leq e}$.
Fix an arbitrary $t\in\mathbb{G}^{n}$ and condition on $\vec{T}=t$ (and on $\mathcal{S}^{\*}$’s internal randomness, so that $\mathcal{S}^{\*}$ is deterministic given the verifier’s replies). By Lemma 1, under this conditioning the secret scalar $r$ is uniform over $\mathbb{F}_{q}$.
We define a sequence of candidate sets $R_{j}\subseteq\mathbb{F}_{q}$ of possible values of $r$ consistent with the transcript up to execution $j$, conditioned on no incorrect acceptance so far. Set $R_{0}:=\mathbb{F}_{q}$.
Consider execution $j\in[e]$ and condition on $\neg\mathsf{Bad}_{\leq j-1}$. Let $\vec{x}^{(j)}$ be the query chosen by $\mathcal{S}^{\*}$ and let $(A^{(j)},B^{(j)})$ be its response. Write $A^{\star(j)}:=\mathsf{MSM}(\vec{P},\vec{x}^{(j)})$.
Case 1: $A^{(j)}=A^{\star(j)}$. Then the verifier’s check $B^{(j)}\stackrel{{\scriptstyle?}}{{=}}rA^{(j)}+\langle\vec{x}^{(j)},\vec{\rho}\rangle Q$ is equivalent to $B^{(j)}\stackrel{{\scriptstyle?}}{{=}}\mathsf{MSM}(\vec{T},\vec{x}^{(j)})=\mathsf{MSM}(t,\vec{x}^{(j)})$, which is independent of $r$ given $\vec{T}=t$. In particular, this execution cannot be an incorrect acceptance, and the verifier’s bit does not shrink the set of possible $r$ values. Thus we set $R_{j}:=R_{j-1}$.
Case 2: $A^{(j)}\neq A^{\star(j)}$. By Lemma 2, for the fixed tuple $(\vec{P},t,\vec{x}^{(j)},A^{(j)},B^{(j)})$ there exists at most one scalar $\hat{r}_{j}\in\mathbb{F}_{q}$ such that the verifier would accept in execution $j$. Therefore, conditioned on $\neg\mathsf{Bad}_{\leq j-1}$ and $\vec{T}=t$, an incorrect acceptance in execution $j$ can occur only if $r=\hat{r}_{j}$; since $r$ is uniform over $R_{j-1}$ under this conditioning, we have
$\Pr[\text{execution $j$ is an incorrect acceptance}\mid\neg\mathsf{Bad}_{\leq j-1},\,\vec{T}=t]\;\leq\;\frac{1}{|R_{j-1}|}.$
Moreover, on the event $\neg\mathsf{Bad}_{\leq j}$ the verifier must have rejected in execution $j$, which implies $r\neq\hat{r}_{j}$ and hence removes at most one candidate value from $R_{j-1}$. Thus we set $R_{j}:=R_{j-1}\setminus\{\hat{r}_{j}\}$ (if $\hat{r}_{j}\in R_{j-1}$), so $|R_{j}|\geq|R_{j-1}|-1$.
Bounding the total failure probability. By construction, $|R_{0}|=q$ and in each execution we remove at most one value of $r$ as long as no incorrect acceptance occurs. Hence on the event $\neg\mathsf{Bad}_{\leq j-1}$ we have $|R_{j-1}|\geq q-(j-1)$, and therefore
$\Pr[\text{execution $j$ is an incorrect acceptance}\mid\neg\mathsf{Bad}_{\leq j-1},\,\vec{T}=t]\;\leq\;\frac{1}{q-(j-1)}.$
Let $p_{j}:=\Pr[\mathsf{Bad}_{\leq j}\mid\vec{T}=t]$. Then
$p_{j}\;\leq\;p_{j-1}+(1-p_{j-1})\cdot\frac{1}{q-(j-1)}.$
Using $p_{0}=0$, a straightforward induction gives $p_{j}\leq j/q$ for all $j\leq e$, and in particular $p_{e}\leq e/q$. Since $t$ was arbitrary, the same bound holds without conditioning on $\vec{T}$, proving the theorem. ∎
5 2G2T: Performance
This section summarizes the concrete performance costs of 2G2T (Algorithm 1) in terms of (i) communication, (ii) server computation, and (iii) client computation. We emphasize the online costs incurred per outsourced MSM query, and separately note the one-time precomputation overhead.
Communication.
2G2T adds only minimal communication beyond what is intrinsically required to outsource an MSM. An untrusted server cannot compute $\mathsf{MSM}(\vec{P},\vec{x})$ without access to the bases $\vec{P}$ and the scalars $\vec{x}$; thus, transmitting $\vec{P}$ (if not already known) and $\vec{x}$ (if not already known) is unavoidable in the generic outsourcing setting.
  • One-time (setup) communication. The client uploads the merged-bases vector $\vec{T}\in\mathbb{G}^{n}$ to the server once, and the server may cache it. If the server does not already have the bases $\vec{P}$, the client also uploads $\vec{P}$ once. Thus the one-time upload is $\vec{T}$ (and optionally $\vec{P}$), i.e., $n$ (or $2n$) group elements.
  • Per-query (online) communication. The server returns exactly two group elements $(A,B)\in\mathbb{G}^{2}$. The client-to-server payload is simply the coefficient vector $\vec{x}$ unless the server already holds it. In particular, in architectures where the server is also the component that produced (or already possesses) the coefficients $\vec{x}$ (e.g., a proving service or an internal accelerator invoked by a prover/verifier), the online communication from client to server can be reduced to a lightweight request/identifier, while the server-to-client response remains the constant-size pair $(A,B)$.
  • Overhead vs. non-verifiable outsourcing. Compared to naïvely outsourcing only $A=\mathsf{MSM}(\vec{P},\vec{x})$, 2G2T adds (i) a one-time upload of $\vec{T}$ and (ii) one additional returned group element $B$ per query.
Server computation.
For each query, the server computes two MSMs:
$A=\mathsf{MSM}(\vec{P},\vec{x})\qquad\text{and}\qquad B=\mathsf{MSM}(\vec{T},\vec{x}).$
Thus, online server work is essentially one additional MSM relative to computing $A$ alone. In practice, implementations based on windowing (e.g., Pippenger-style bucket methods) can reuse the scalar decomposition of $\vec{x}$ across both MSMs, since the scalars are identical and only the bases differ. Consequently, the incremental cost of the second MSM is typically less than a full second pass of all scalar preprocessing.
Client computation.
Per query, the client performs no large MSM in $\mathbb{G}$. Instead, verification consists of:
  • one length-$n$ inner product in the scalar field,
    $s=\langle\vec{x},\vec{\rho}\rangle=\sum_{i=1}^{n}x_{i}\rho_{i}\in\mathbb{F}_{q},$
  • two scalar multiplications in $\mathbb{G}$ (compute $rA$ and $sQ$),
  • one group addition (compute $rA+sQ$) and one equality check against $B$.
In particular, the number of group operations performed by the client per query is constant, independent of $n$. As $n$ grows, the dominant client cost is the field-arithmetic inner product, while the elliptic-curve/group work remains essentially fixed.
One-time precomputation overhead.
The precomputation phase computes $\vec{T}=(T_{1},\dots,T_{n})$ with $T_{i}=rP_{i}+\rho_{i}Q$ for all $i$. This incurs $n$ scalar multiplications by a fixed scalar $r$ on the bases $\vec{P}$, $n$ scalar multiplications on the fixed base $Q$, and $n$ group additions. Because this phase is executed once and amortized over many outsourced queries, its cost is typically dominated by the online savings when many MSMs are outsourced over the same bases $\vec{P}$. Moreover, fixed-base techniques can accelerate the $Q$-multiplications, since $Q$ is constant.
6 Related work
Verifiable MSM in transparent, pairing-free groups.
The closest related line of work that explicitly targets verifiable multi-exponentiation/MSM in transparent groups is due to Pinkas [18], which gives a non-interactive, publicly verifiable statistical check with soundness parameter $2^{-\lambda}$. At a high level, Pinkas’ protocol attains public verifiability by returning a proof consisting of on the order of $\log q$ group elements and by performing verifier work that depends on the target soundness parameter $\lambda$. In contrast, 2G2T is a designated-verifier protocol with a one-time keyed setup whose per-query verification is dominated by a single length-$n$ field inner product plus $O(1)$ group operations, while achieving constant-size server-to-client communication (two group elements) and per-query soundness error $1/q$ in prime-order groups. We view these as complementary points in the design space: public verifiability with larger proofs versus designated-verifier verification with constant-size responses and “full-group” statistical soundness.
A second line of work considers MSM outsourcing for zero-knowledge proof generation. In particular, Xiong et al. [21] propose an “immediate verification” step based on a consistency relation between multiple values returned by the server. However, as stated, such a consistency check does not by itself force the server to incorporate all $n$ terms of the MSM instance: a malicious server can satisfy the stated relation while computing only a strict subset of indices (e.g., returning a partial sum such as $A=x_{1}P_{1}$), even though $A\neq\mathsf{MSM}(\vec{P},\vec{x})$ in general. Thus, in the standard adversarial-server model for outsourcing, the described immediate-check mechanism does not, on its own, imply soundness.
Structured (non-transparent) approaches via SRS and pairings.
In bilinear (pairing-based) settings with a structured reference string (SRS), many large algebraic relations can be certified succinctly via a constant number of pairing equations. A canonical example is the Kate–Zaverucha–Goldberg (KZG) polynomial commitment scheme: the commitment is an MSM against SRS-derived bases, yet a constant-size evaluation proof can be verified using $O(1)$ pairings, independent of the polynomial degree [14]. This paradigm underlies many non-transparent SNARK constructions with very fast verification (e.g., Groth16 and polynomial-commitment–based families), where the proof system itself provides an integrity layer that certifies the relevant algebraic relations without requiring the verifier to redo large MSMs [13, 15, 6]. Our work targets a different regime: transparent, pairing-free prime-order groups and generic-base MSM instances, where the bases are not SRS-derived and there is no pairing-based shortcut for succinct certification.
Homomorphic authenticators and authentication-based delegation.
Our check is closely related in spirit to homomorphic authenticators (a.k.a. homomorphic MACs/signatures) for linear functions: a client authenticates a dataset once, an untrusted server computes on authenticated data, and the verifier checks that the output is the correct function evaluation. Linearly homomorphic MACs were introduced in the context of network coding, where intermediate nodes forward linear combinations of packets while preserving integrity [1], and subsequent work generalized homomorphic authentication to broader classes of computations and more practical instantiations [11, 5]. 2G2T can be viewed as an especially simple linearly-homomorphic authenticator specialized to MSM: the public merged-bases vector $\vec{T}$ plays the role of an evaluation key, while the verifier’s secret state binds the server to the entire MSM instance.
General verifiable computation.
Finally, MSM outsourcing can be viewed as a special case of verifiable computation and delegation, where the goal is to verify correctness with substantially less work than direct evaluation. General-purpose frameworks such as GKR-style interactive proofs and non-interactive verifiable computing formalize this goal, but are typically heavier than what is desirable for a ubiquitous cryptographic kernel [12, 9].
7 Performance evaluation
This section empirically evaluates the computational cost of 2G2T and quantifies the speedup obtained by verifying an outsourced MSM (via the 2G2T check) rather than computing the MSM locally. All benchmarks were run on an Apple M4 machine with 32 GB of memory running macOS Tahoe, compiled with rustc 1.93.0 (254b59607 2026-01-19) and using curve25519-dalek v4.1.3. Figure 1 summarizes our main results.
Implementation setting and curve choice.
We implement a reference benchmark in Rust using the curve25519-dalek library over the Ristretto255 group. Ristretto255 is a natural instantiation for our target setting: it provides a prime-order group abstraction, it is transparent (no pairings, no SRS), and it is commonly used in Bulletproofs-style deployments. Throughout, we fix the public point $Q$ to the standard Ristretto basepoint so that the verifier’s fixed-base multiplication $sQ$ can leverage the library’s precomputed basepoint table. [^2] The group-only part of verification is already negligible in our measurements; choosing $Q$ as the basepoint simply reflects a realistic engineering choice for an always-fixed public $Q$.
For each dimension $n=2^{e}$ (with $e\in\{10,\dots,18\}$), we sample $\vec{P}\in\mathbb{G}^{n}$ and $\vec{x}\in\mathbb{F}_{q}^{n}$ at random, run the 2G2T precomputation once to obtain $\vec{T}$, and then measure: (i) the time to compute $\mathsf{MSM}(\vec{P},\vec{x})$ under selected baselines, and (ii) the time for the client-side 2G2T verification (i.e., compute $s=\langle\vec{x},\vec{\rho}\rangle$ and check $B\stackrel{{\scriptstyle?}}{{=}}rA+sQ$). We report per-iteration averages over many iterations (after warmup) and focus on ratios (speedups), which are relatively stable across machines.
Verification cost and “lazy” inner product.
In 2G2T, the verifier’s dominant term is computing the scalar
$s=\langle\vec{x},\vec{\rho}\rangle=\sum_{i=1}^{n}x_{i}\rho_{i}\in\mathbb{F}_{q},$
followed by a constant amount of group work ($rA$, sQsQ, one addition, one comparison). To reduce modular-reduction overhead in the scalar field, we implement a lazy-reduction inner product: we accumulate products in a wide (512-bit) accumulator and reduce modulo $q$ only once per block of terms. In the experiments of Figure 1, we use a block size of $128$, which keeps the wide accumulator safely within 512 bits while substantially reducing the number of reductions. This optimization directly targets the verifier bottleneck and is faithful to how a high-performance 2G2T verifier would be engineered.
Baselines and why we include a “naïve” MSM.
Figure 1 plots speedup curves against two MSM baselines:
  • Optimized MSM baseline (library MSM). We use curve25519-dalek’s optimized multi-scalar multiplication routine
    (RistrettoPoint::vartime_multiscalar_mul). This is the appropriate baseline for performance-oriented deployments and captures the best available software MSM in our chosen setting.
  • Naïve MSM baseline (sum of scalar multiplications). We also measure the straightforward method
    $\sum_{i=1}^{n}x_{i}P_{i}\quad\text{computed as}\quad\text{acc}\leftarrow 0;\ \text{for }i:\ \text{acc}\leftarrow\text{acc}+x_{i}P_{i}.$
    While asymptotically suboptimal, this baseline is still meaningful in practice and is worth reporting: (i) it approximates behavior on memory-constrained devices where bucket-based MSM is unattractive due to RAM/cache pressure; (ii) it reflects implementations that prioritize simplicity/auditability or constant-time scalar multiplication primitives (where MSM optimizations are not available or are disabled); and (iii) it provides a clear reference point that isolates how much of the gain comes from 2G2T verification versus advanced MSM engineering.
We emphasize that the optimized MSM baseline is the primary one; the naïve MSM curve is included as an additional, practically motivated reference.
Results and interpretation.
Figure 1 reports the speedup defined as
$\mathsf{Speedup}(n)\;:=\;\frac{t_{\mathrm{MSM}}(n)}{t_{\mathrm{verify}}(n)},$
where $t_{\mathrm{MSM}}(n)$ is the time to compute $\mathsf{MSM}(\vec{P},\vec{x})$ under the chosen baseline and $t_{\mathrm{verify}}(n)$ is the client-side 2G2T verification time.
The key observation is that 2G2T verification is extremely lightweight even for large $n$: the group-only part of the check is essentially constant (two scalar multiplications with one fixed base, one addition, one equality test), while the only $n$-dependent work is the inner product. Consequently, the speedup curve increases with $n$ (as fixed verification overhead is amortized) and then plateaus once verification is dominated by the linear-time inner product. At the largest tested dimension ($n=2^{18}=262{,}144$), verification takes on the order of a few milliseconds in our implementation, whereas recomputing the MSM requires hundreds of milliseconds under the optimized MSM routine and several seconds under the naïve baseline. This corresponds to speedups on the order of hundreds of times against the optimized MSM baseline and thousands of times against the naïve baseline at large $n$.
Overall, the measurements support the main performance claim of 2G2T: the verifier can check an outsourced MSM with (i) constant-size server response (two group elements), (ii) essentially constant group work, and (iii) only a single streaming pass over $\vec{x}$ for the inner product, yielding large end-to-end speedups in the large-$n$ regime relevant to aggregated/inner-product-based proof systems.
Figure 1: 2G2T speedup on Ristretto255 (up to $\sim 300\times$ vs. optimized MSM and $\sim 3000\times$ vs. naïve MSM, for $n\leq 2^{18}$).
Comparison to Pinkas.
For context, Pinkas [18] proposes a publicly verifiable statistical check for MSMs in prime-order groups. In his analysis, when scalars have bit-length $m$ (typically $m\approx\log_{2}q$), verification reduces to computing an $n$-wide MSM with shorter $(\lambda+\log_{2}m)$-bit exponents, yielding an expected speedup of roughly $\frac{m}{\lambda+\log_{2}m}$. Over Ristretto255, one may take $m\approx\log_{2}q\approx 252$ and $\log_{2}m\approx 8$, which predicts a speedup of about $252/(64+8)\approx 3.5\times$ at $\lambda=64$. In contrast, matching 2G2T’s per-query soundness error $1/q$ corresponds to $\lambda\approx\log_{2}q$ (about $252$), in which case the predicted speedup essentially vanishes. Pinkas’ protocol also requires the server to return $m$ intermediate group elements (about 252252256256 for Ristretto255), whereas 2G2T returns only two group elements $(A,B)$. These protocols therefore occupy different points in the design space: Pinkas provides public verifiability with $\Theta(\log q)$ communication, while 2G2T achieves constant-size responses and very fast designated-verifier checks after a one-time keyed setup.
8 Conclusion
We introduced 2G2T, a lightweight protocol for verifiably outsourcing generic-base multi-scalar multiplication $\mathsf{MSM}(\vec{P},\vec{x})$ in transparent prime-order groups. Our analysis shows that 2G2T is perfectly complete and achieves full-group statistical soundness: even against an unbounded adversarial server, the probability of accepting an incorrect $A$ is at most $1/q$ per execution, and at most $e/q$ over $e$ adaptive executions, requiring only that $\mathbb{G}$ has prime order $q$. Empirically, a Rust implementation over Ristretto255 confirms that verification is extremely fast in the large-$n$ regime: for $n$ up to $2^{18}$, verification is up to $\sim 300\times$ faster than locally computing an optimized MSM, while the server-to-client response remains constant-size (two compressed group elements, 64 bytes on Ristretto255). Overall, 2G2T isolates a simple and practical primitive for integrity-preserving MSM acceleration with minimal communication and verifier overhead.
Limitations and future work.
2G2T is a designated-verifier protocol with a one-time keyed setup for fixed bases, and we do not aim to hide the coefficients $\vec{x}$ from the server. An important direction is extending 2G2T to prover-side outsourcing, where $\vec{x}$ may be witness-dependent and must remain hidden, while preserving verification that is far cheaper than a full MSM. A second direction is exploring publicly verifiable variants, likely trading off larger proofs and/or stronger assumptions for removing the verifier’s secret state.

### v2

@@
-arXiv:2602.23464v1[cs.CR] 26 Feb 2026
+2G2T: Constant-Size, Statistically Sound MSM Outsourcing
 Abstract
@@
 We present 2G2T, a simple protocol for verifiably outsourcing MSM to an untrusted server.
-After a one-time keyed setup for fixed bases $\vec{P}$ (producing a public merged-bases vector $\vec{T}$ and client secret state), the server returns only two group elements per query: the claimed MSM result $A=\mathsf{MSM}(\vec{P},\vec{x})$ and an auxiliary value $B=\mathsf{MSM}(\vec{T},\vec{x})$.
-Verification requires a single length-$n$ field inner product and a constant number of group operations (two scalar multiplications and one addition), while the server performs two MSMs.
-In our Ristretto255 implementation, verification is up to $\sim 300\times$ faster than computing the MSM locally using a highly optimized MSM routine for $n$ up to $2^{18}$, and the server-to-client response is constant-size (two compressed group elements, 64 bytes on Ristretto255).
-Despite its simplicity and efficiency, 2G2T achieves statistical soundness: for any (even unbounded) adversarial server, the probability of accepting an incorrect result is at most $1/q$ per query, and at most $e/q$ over $e$ adaptive executions, in a prime-order group of size $q$.
+2G2T is efficient for both parties: the server performs only two MSM computations and returns only two group elements to the client, namely the claimed result $A=\mathsf{MSM}(\vec{P},\vec{x})$ and an auxiliary group element $B$.
+Client-side verification consists of a single length-$n$ field inner product and only three group operations (two scalar multiplications and one group addition).
+In our Ristretto255 implementation, verification is up to $\sim 300\times$ faster than computing the MSM locally using a highly optimized MSM routine (for $n$ up to $2^{18}$).
+Moreover, 2G2T enables latency-hiding verification: nearly all verifier work can be performed while waiting for the server’s response, so once $(A,B)$ arrives the verifier completes the check with only one scalar multiplication and one group addition (both independent of $n$).
+Finally, despite its simplicity and efficiency, we prove that 2G2T achieves statistical soundness: for any (even unbounded) adversarial server, the probability of accepting an incorrect result is at most $1/q$ per query, and at most $e/q$ over $e$ adaptive executions, in a prime-order group of size $q$.
 1 Introduction
@@
 This motivates specialized protocols tailored to MSM that provide strong integrity guarantees with minimal client-side verification overhead.
-Challenges.
-Prior work on verifiable MSM outsourcing highlights two obstacles: (i) achieving a meaningful speedup while keeping statistical soundness extremely high, and (ii) avoiding lightweight checks that enforce only a consistency relation between returned values but do not bind the server to all $n$ terms of the instance.
-For example, Pinkas [18] gives a non-interactive, publicly verifiable statistical check that can yield a modest constant-factor speedup when one targets moderate soundness (on the order of tens of bits), but doing so typically increases the server’s response size to hundreds of group elements on 255-bit groups, and the speedup diminishes as the soundness target approaches “full-group” levels.
-At the other end of the spectrum, cheaper consistency-only strategies can fail outright against a malicious server: a server may compute an MSM over only a strict subset of indices (e.g., return $A=x_{1}P_{1}$) and still satisfy the verifier’s consistency relation, even though $A\neq\mathsf{MSM}(\vec{P},\vec{x})$ (see, e.g., [21]).
-Together, these examples underscore that verifiable MSM outsourcing is nontrivial: one must simultaneously achieve negligible soundness error and bind the server to the full MSM instance without reintroducing MSM-like verifier work.
-Our contribution: 2G2T.
+Our proposed method: 2G2T.
 We present 2G2T, a statistically sound protocol for verifiable MSM outsourcing with constant-size server-to-client communication and fast client-side verification.
-[^1] The mnemonic 2G2T reflects (i) “two group elements to transfer”: the server returns two group elements per query; and (ii) it can seem “too good to be true” at first glance, as it combines constant-size communication, $O(1)$ verifier group operations, and $1/q$-level statistical soundness.
-After a one-time setup for fixed bases $\vec{P}$, the server answers each client’s query $\vec{x}$ with only two group elements: $A$ claimed to equal $\mathsf{MSM}(\vec{P},\vec{x})$ and an auxiliary value $B=\mathsf{MSM}(\vec{T},\vec{x})$ for a precomputed merged-bases vector $\vec{T}$.
-The client verifies using a single length-$n$ field inner product $s=\langle\vec{x},\vec{\rho}\rangle$ and a constant number of group operations (two scalar multiplications and one addition), while the server performs exactly two MSMs (only one more than computing $A$ alone).
-We prove that 2G2T is perfectly complete and statistically sound: for any (even computationally unbounded) adversarial server, the probability that the verifier accepts an incorrect $A$ is at most $1/q$ per execution, and at most $e/q$ over $e$ adaptive executions (Theorems 4.14.3).
-Our Rust implementation over Ristretto255 confirms that 2G2T yields large practical gains in the large-$n$ regime: verification is up to $\sim 300\times$ faster than computing a highly optimized MSM, and up to $\sim 3{,}000\times$ faster than a naïve “$n$ scalar multiplications” baseline, for dimensions up to $n=2^{18}=262{,}144$ (with exact constants depending on hardware and implementation details).
-Finally, server-to-client communication is constant and minimal: the response consists of two compressed group elements $(A,B)$, i.e., 64 bytes in Ristretto255, independent of $n$.
-These efficiency gains come while providing “full-group” statistical integrity: the per-query soundness error $1/q$ corresponds to about $\log_{2}(q)$ bits of soundness (roughly $252$ bits for Ristretto255).
+[^1] The mnemonic 2G2T reflects (i) “two group elements to transfer”: the server returns two group elements per query; and (ii) that the protocol can seem “too good to be true” at first glance, as it combines constant-size communication, $O(1)$ verifier group operations, and $1/q$-level statistical soundness.
+After a one-time setup for fixed bases $\vec{P}$, the server answers a query $\vec{x}$ with only two group elements: AA claimed to equal $\mathsf{MSM}(\vec{P},\vec{x})$ and an auxiliary value $B=\mathsf{MSM}(\vec{T},\vec{x})$ for a precomputed public merged-bases vector $\vec{T}$.
+The client verifies using a single length-$n$ field inner product and $O(1)$ group operations (two scalar multiplications and one addition), while the server performs exactly two MSMs (only one more than computing AA alone).
+Our Rust implementation over Ristretto255 confirms that 2G2T yields large practical gains in the large-$n$ regime: verification is up to $\sim 300\times$ faster than computing a highly optimized MSM, and up to $\sim 3{,}000\times$ faster than a naïve “$n$ scalar multiplications” baseline, for dimensions up to $n=2^{18}$.
+Moreover, verification is latency-hiding: essentially all verifier work (the inner product and one fixed-base scalar multiplication) can be performed while waiting for the server’s response, so once $(A,B)$ arrives the client performs only one scalar multiplication and one group addition to complete the check.
+We prove that 2G2T is perfectly complete and statistically sound: for any (even computationally unbounded) adversarial server, the probability that the verifier accepts an incorrect AA is at most $1/q$ per execution (roughly 252252 bits for Ristretto255), and at most $e/q$ over $e$ adaptive executions (Theorems 4.14.3).
+Encrypted MSM (EMSM) and composability.
+The closest prior work that directly targets MSM outsourcing against an untrusted server while also hiding the scalar vector $\vec{x}$ is the recent EMSM line of work [1], motivated primarily by prover-side outsourcing where $\vec{x}$ is sensitive and must remain hidden from the server.
+In contrast, we focus on the common verifier-side regime where $\vec{x}$ is public (or already known to the server) and the bottleneck is simply evaluating large MSMs.
+This includes, for example, SNARK/Bulletproofs-style verification, where the MSM scalars are derived from the public transcript and public inputs, and deployments where the prover already holds $\vec{x}$ and can act as the server.
+In this public-scalar setting, 2G2T provides an extremely lightweight integrity layer: the client verifies using a single length-$n$ inner product plus $O(1)$ curve operations, and verification is naturally latency-hiding.
+[^2] As reported in [1], EMSM’s client time for large instances is dominated by a pre-query encryption step that must complete before the request can be sent, which leaves substantially less opportunity for latency-hiding.
+When scalar privacy is required, 2G2T composes cleanly with EMSM: EMSM can provide coefficient hiding, while 2G2T supplies the integrity check.
+This avoids the additional EMSM execution used to harden against an untrusted server, reducing client-side work by up to a factor of two, without increasing the server’s MSM work or the number of group elements returned (relative to EMSM’s untrusted server setting).
+At the same time, the integrity guarantee is upgraded to statistical soundness, rather than relying on EMSM’s dual-LPN-style assumptions.
 2 Threat model and security goals
@@
 An MSM outsourcing protocol should satisfy the following standard properties.
-  • Completeness. If $\mathcal{S}$ follows the protocol and returns correct values, then $\mathcal{C}$ accepts and outputs $Y=\mathsf{MSM}(\vec{P},\vec{x})$ for every query $\vec{x}$.
+  • Completeness. If $\mathcal{S}$ follows the protocol and returns correct values, then $\mathcal{C}$ accepts and outputs $A=\mathsf{MSM}(\vec{P},\vec{x})$ for every query $\vec{x}$.
   • Soundness (integrity).
@@
 $\Pr[r=a\mid\vec{T}=t]\;=\;\Pr[r=a]\;=\;\frac{1}{q}.$
-Lemma 2(Uniqueness of an accepting scalar for an incorrect $A$)
+Lemma 2(Uniqueness of an accepting scalar for an incorrect AA)
 Fix $\vec{P}\in\mathbb{G}^{n}$, $\vec{T}\in\mathbb{G}^{n}$, a query $\vec{x}\in\mathbb{F}_{q}^{n}$, and a pair $(A,B)\in\mathbb{G}^{2}$.
@@
 $\Pr\Bigl[\,\mathsf{Vrfy}(r,\vec{\rho};\vec{x},A,B)=1\ \wedge\ A\neq\mathsf{MSM}(\vec{P},\vec{x})\,\Bigr]\;\leq\;\frac{1}{q},$
-where the probability is over the client’s randomness in the precomputation of Definition 1.
+where the probability is over the client’s randomness in the setup and any internal randomness used by the (possibly randomized) server.
 Proof
@@
 $A=\mathsf{MSM}(\vec{P},\vec{x})\qquad\text{and}\qquad B=\mathsf{MSM}(\vec{T},\vec{x}).$
-Thus, online server work is essentially one additional MSM relative to computing $A$ alone.
+Thus, online server work is essentially one additional MSM relative to computing AA alone.
 In practice, implementations based on windowing (e.g., Pippenger-style bucket methods) can reuse the scalar decomposition of $\vec{x}$ across both MSMs, since the scalars are identical and only the bases differ.
@@
     $s=\langle\vec{x},\vec{\rho}\rangle=\sum_{i=1}^{n}x_{i}\rho_{i}\in\mathbb{F}_{q},$
-  • two scalar multiplications in $\mathbb{G}$ (compute $rA$ and $sQ$),
+  • two scalar multiplications in $\mathbb{G}$ (compute rArA and sQsQ),
   • one group addition (compute $rA+sQ$) and one equality check against $B$.
@@
 One-time precomputation overhead.
-The precomputation phase computes $\vec{T}=(T_{1},\dots,T_{n})$ with $T_{i}=rP_{i}+\rho_{i}Q$ for all $i$.
+The precomputation phase computes $\vec{T}=(T_{1},\dots,T_{n})$ with $T_{i}=rP_{i}+\rho_{i}Q$ for all ii.
 This incurs $n$ scalar multiplications by a fixed scalar $r$ on the bases $\vec{P}$, $n$ scalar multiplications on the fixed base $Q$, and $n$ group additions.
@@
 The closest related line of work that explicitly targets verifiable multi-exponentiation/MSM in transparent groups is due to Pinkas [19], which gives a non-interactive, publicly verifiable statistical check with soundness parameter $2^{-\lambda}$.
-At a high level, Pinkas’ protocol attains public verifiability by returning a proof consisting of on the order of $\log q$ group elements and by performing verifier work that depends on the target soundness parameter $\lambda$.
+At a high level, Pinkas’ protocol attains public verifiability by returning a proof consisting of $O(\log q)$ group elements and by performing verifier work that depends on the target soundness parameter $\lambda$.
 In contrast, 2G2T is a designated-verifier protocol with a one-time keyed setup whose per-query verification is dominated by a single length-$n$ field inner product plus $O(1)$ group operations, while achieving constant-size server-to-client communication (two group elements) and per-query soundness error $1/q$ in prime-order groups.
@@
 [22] propose an “immediate verification” step based on a consistency relation between multiple values returned by the server.
-However, as stated, such a consistency check does not by itself force the server to incorporate all $n$ terms of the MSM instance: a malicious server can satisfy the stated relation while computing only a strict subset of indices (e.g., returning a partial sum such as $A=x_{1}P_{1}$), even though $A\neq\mathsf{MSM}(\vec{P},\vec{x})$ in general.
+However, such a consistency check does not by itself force the server to incorporate all $n$ terms of the MSM instance: a malicious server can satisfy the stated relation while computing only a strict subset of indices (e.g., returning a partial sum such as $A=x_{1}P_{1}$), even though $A\neq\mathsf{MSM}(\vec{P},\vec{x})$ in general.
 Thus, in the standard adversarial-server model for outsourcing, the described immediate-check mechanism does not, on its own, imply soundness.
@@
 Ristretto255 is a natural instantiation for our target setting: it provides a prime-order group abstraction, it is transparent (no pairings, no SRS), and it is commonly used in Bulletproofs-style deployments.
-Throughout, we fix the public point $Q$ to the standard Ristretto basepoint so that the verifier’s fixed-base multiplication $sQ$ can leverage the library’s precomputed basepoint table.
+Throughout, we fix the public point $Q$ to the standard Ristretto basepoint so that the verifier’s fixed-base multiplication sQsQ can leverage the library’s precomputed basepoint table.
 [^3] The group-only part of verification is already negligible in our measurements; choosing $Q$ as the basepoint simply reflects a realistic engineering choice for an always-fixed public $Q$.
@@
 $s=\langle\vec{x},\vec{\rho}\rangle=\sum_{i=1}^{n}x_{i}\rho_{i}\in\mathbb{F}_{q},$
-followed by a constant amount of group work ($rA$, sQsQ, one addition, one comparison).
+followed by a constant amount of group work (rArA, sQsQ, one addition, one comparison).
 To reduce modular-reduction overhead in the scalar field, we implement a lazy-reduction inner product: we accumulate products in a wide (512-bit) accumulator and reduce modulo $q$ only once per block of terms.
-In the experiments of Figure 1, we use a block size of $128$, which keeps the wide accumulator safely within 512 bits while substantially reducing the number of reductions.
+In the experiments of Figure 1, we use a block size of 128128, which keeps the wide accumulator safely within 512 bits while substantially reducing the number of reductions.
 This optimization directly targets the verifier bottleneck and is faithful to how a high-performance 2G2T verifier would be engineered.
@@
 Over Ristretto255, one may take $m\approx\log_{2}q\approx 252$ and $\log_{2}m\approx 8$, which predicts a speedup of about $252/(64+8)\approx 3.5\times$ at $\lambda=64$.
-In contrast, matching 2G2T’s per-query soundness error $1/q$ corresponds to $\lambda\approx\log_{2}q$ (about $252$), in which case the predicted speedup essentially vanishes.
+In contrast, matching 2G2T’s per-query soundness error $1/q$ corresponds to $\lambda\approx\log_{2}q$ (about 252252), in which case the predicted speedup essentially vanishes.
 Pinkas’ protocol also requires the server to return $m$ intermediate group elements (about 252252256256 for Ristretto255), whereas 2G2T returns only two group elements $(A,B)$.
@@
 We introduced 2G2T, a lightweight protocol for verifiably outsourcing generic-base multi-scalar multiplication $\mathsf{MSM}(\vec{P},\vec{x})$ in transparent prime-order groups.
-Our analysis shows that 2G2T is perfectly complete and achieves full-group statistical soundness: even against an unbounded adversarial server, the probability of accepting an incorrect $A$ is at most $1/q$ per execution, and at most $e/q$ over $e$ adaptive executions, requiring only that $\mathbb{G}$ has prime order $q$.
+Our analysis shows that 2G2T is perfectly complete and achieves full-group statistical soundness: even against an unbounded adversarial server, the probability of accepting an incorrect AA is at most $1/q$ per execution, and at most $e/q$ over $e$ adaptive executions, requiring only that $\mathbb{G}$ has prime order $q$.
 Empirically, a Rust implementation over Ristretto255 confirms that verification is extremely fast in the large-$n$ regime: for $n$ up to $2^{18}$, verification is up to $\sim 300\times$ faster than locally computing an optimized MSM, while the server-to-client response remains constant-size (two compressed group elements, 64 bytes on Ristretto255).